# FullLoRA-AT: Efficiently Boosting the Robustness of Pretrained Vision Transformers

### v1

@@ 4.3 Comparison with Other Methods
FullLoRA-AT (ours, r=16) 2.46 62.34 24.99 26.31 26.19 23.69 8.00
FullLoRA-AT (ours, r=32) 4.85 61.20 25.86 28.40 28.23 25.78 8.38
Imagenette Fully Finetune 85.15 93.40 69.00 68.80 68.60 67.60 8.52
Patch Embedding Layer 0.04 37.60 23.40 25.60 25.60 24.70 5.76
Layer Normalization 0.04 61.40 23.60 28.80 28.80 27.60 5.89
Aurora [52] 0.09 87.10 51.30 53.20 52.90 51.80 9.07
LoRA (r=32) [20] 4.72 94.00 63.80 62.00 61.20 60.00 6.15
UniPELT [35] 11.42 91.40 58.00 57.80 56.20 55.30 9.58
FullLoRA-AT (ours, r=16) 2.46 93.70 64.10 63.70 63.00 62.10 6.01
FullLoRA-AT (ours, r=32) 4.85 93.40 66.20 66.20 65.60 64.00 6.32
Table 3: Comparison with other methods on different ViT-based models.
Model Method Trainable Clean Acc CW-20 PGD-20 PGD-100 AutoAttack Time
Paras (M) (%) (%) (%) (%) (%) (h)
ViT-B Fully Finetune 85.15 82.83 50.26 52.85 52.56 48.26 11.17
Aurora [52] 0.09 76.09 38.50 41.85 41.67 39.87 11.48
LoRA (r=32) [20] 4.72 84.11 47.14 49.96 49.74 45.13 8.10
UniPELT [35] 11.42 84.63 47.63 49.49 48.99 44.63 13.11
FullLoRA-AT (ours, r=16) 2.46 85.93 47.68 50.33 49.46 45.82 8.01
FullLoRA-AT (ours, r=32) 4.85 87.12 49.76 51.38 50.83 46.94 8.38
ViT-S Fully Finetune 47.34 79.94 48.02 50.84 50.62 46.22 5.31
Aurora [52] 0.08 70.52 36.12 38.91 38.79 35.24 5.46
LoRA (r=32) [20] 3.15 78.31 44.87 46.10 45.98 41.37 3.87
UniPELT [35] 7.75 73.40 39.82 43.68 43.49 40.09 5.93
FullLoRA-AT (ours, r=16) 1.65 78.66 45.10 46.93 46.60 42.31 3.69
FullLoRA-AT (ours, r=32) 3.25 79.03 45.63 48.58 48.42 44.93 3.92
Swin-B Fully Finetune 86.69 81.29 48.36 50.99 50.88 46.57 6.44
Aurora [52] 0.09 75.13 37.52 38.17 38.03 35.26 4.79
LoRA (r=32) [20] 6.16 80.94 45.32 46.50 46.37 43.08 4.66
UniPELT [35] 11.58 79.55 44.16 45.28 44.93 41.39 5.07
FullLoRA-AT (ours, r=16) 3.17 82.86 46.10 47.53 47.34 43.57 4.49
FullLoRA-AT (ours, r=32) 6.29 83.67 46.73 48.16 48.05 44.16 4.81
Table 4: Results of our FullLoRA-AT under different rank $r$.
@@ 4.4 Results under different rank $r$
We compare the robustness of pretrained models adversarially finetuned using our proposed FullLoRA-AT method with different ranks $r$ in the LNLoRA module. As shown in Tab. 4, with the increase of $r$, there is a corresponding increase in the number of trainable parameters during finetuning, as well as a slight increase in the required training duration. Concurrently, the robustness of the resulting model also noticeably improves. When $r=8$, the model finetuned using our FullLoRA-AT method, with only about 1.5% of the trainable parameters used for full finetuning, incurs only a 3%-4% loss in robustness under various attacks. Increasing $r$ to 64, which utilizes 11.3% of the training parameters, further reduces the robustness loss to approximately 1%. Considering the trade-off between the number of parameters used for adversarial finetuning and the robustness of the model, we choose rank $r=32$ as the default for other experiments in our work.
4.5 Combined with other advanced adversarial training methods
Our method can also be further combined with other advanced adversarial training methods to enhance the results of our adversarial finetuning.
In Tab. 5, we present the results of combining FullLoRA-AT with methods designed to accelerate the speed of adversarial training. In our previous experiments, we use a 10-step PGD [33] to generate adversarial examples for training, referred to as PGD-10. When we switch to PGD-5, the time required for adversarial finetuning on pretrained models is reduced, and our method consistently outperforms the LoRA [20] in terms of robustness. Similarly, when combined with the AGAT [57], an attention-guided tokens discarding approach, the adversarial finetuning time is further reduced, albeit at the expense of a slight decrease in model robustness. In summary, when combined with different advanced methods for accelerating adversarial training, our FullLoRA-AT consistently achieves higher model robustness on various datasets compared to the LoRA method, demonstrating its broad applicability.
Additionally, we also combine our FullLoRA-AT with other methods aimed at enhancing the performance of adversarial robustness. As shown in Tab. 6, when combined with MART [54] and PRM [36], our FullLoRA-AT method still significantly outperforms the LoRA method in terms of model robustness across different datasets.
@@ 4.5 Combined with other advanced adversarial training methods
Table 5: Results of combining FullLoRA-AT with other speed-accelerating adversarial training methods.
Dataset Combined Method Trainable Clean Acc CW-20 PGD-20 PGD-100 AutoAttack Time
Method Paras (M) (%) (%) (%) (%) (%) (h)
CIFAR-10 PGD-5 [33] Fully Finetune 85.15 86.98 49.97 51.08 50.52 47.58 6.37
LoRA [20] 4.72 86.10 45.09 45.23 45.01 42.46 4.10
FullLoRA-AT (ours) 4.85 89.30 47.66 47.92 47.18 44.97 4.36
AGAT [57] Fully Finetune 85.15 87.42 45.02 47.75 47.54 42.86 5.76
LoRA [20] 4.72 86.34 42.68 43.89 43.65 39.68 3.81
FullLoRA-AT (ours) 4.85 89.47 44.07 45.00 44.45 41.03 3.93
CIFAR-100 PGD-5 [33] Fully Finetune 85.22 63.27 27.21 27.64 27.57 25.88 6.39
LoRA [20] 4.72 65.16 23.49 24.87 24.63 21.06 4.11
FullLoRA-AT (ours) 4.85 66.49 24.75 26.10 26.01 24.49 4.38
AGAT [57] Fully Finetune 85.22 63.91 25.36 25.74 25.61 24.37 5.77
LoRA [20] 4.72 65.18 21.73 21.99 21.86 20.42 3.82
FullLoRA-AT (ours) 4.85 66.31 22.93 23.07 23.01 22.54 3.93
Imagenette PGD-5 [33] Fully Finetune 85.15 94.40 64.60 64.20 63.60 62.80 4.46
LoRA [20] 4.72 93.10 60.70 58.90 58.30 57.60 3.31
FullLoRA-AT (ours) 4.85 94.00 63.80 62.00 61.20 60.30 3.37
AGAT [57] Fully Finetune 85.15 94.70 61.90 61.60 61.10 60.60 4.15
LoRA [20] 4.72 93.00 57.20 56.20 56.00 55.10 3.19
FullLoRA-AT (ours) 4.85 95.20 60.20 60.00 59.70 58.90 3.23
@@
Table 6: Results of combining FullLoRA-AT with other performance-enhancing adversarial training methods.
Dataset Combined Method Trainable Clean Acc CW-20 PGD-20 PGD-100 AutoAttack Time
Method Paras (M) (%) (%) (%) (%) (%) (h)
CIFAR-10 MART [54] Fully Finetune 85.15 83.65 51.09 53.62 53.41 49.02 12.13
LoRA [20] 4.72 84.82 47.93 50.77 50.55 46.01 8.86
FullLoRA-AT (ours) 4.85 87.99 50.59 52.09 51.64 47.66 8.97
PRM [36] Fully Finetune 85.15 83.43 50.97 53.49 53.12 48.81 11.23
LoRA [20] 4.72 84.62 47.71 50.53 50.45 45.78 8.16
FullLoRA-AT (ours) 4.85 87.76 50.39 51.82 51.42 47.36 8.39
CIFAR-100 MART [54] Fully Finetune 85.22 61.82 29.51 31.46 31.20 28.23 12.15
LoRA [20] 4.72 61.70 25.44 26.59 26.60 23.36 8.88
FullLoRA-AT (ours) 4.85 62.00 26.60 29.22 29.13 26.68 8.99
PRM [36] Fully Finetune 85.22 61.46 29.15 31.12 31.12 28.13 11.25
LoRA [20] 4.72 62.59 25.11 26.26 26.17 22.97 8.18
FullLoRA-AT (ours) 4.85 61.81 26.35 28.88 28.70 26.35 8.40
Imagenette MART [54] Fully Finetune 85.15 94.26 69.81 69.68 69.40 68.49 9.32
LoRA [20] 4.72 94.06 64.57 62.78 62.00 60.88 6.75
FullLoRA-AT (ours) 4.85 94.19 66.99 67.10 66.51 64.91 6.83
PRM [36] Fully Finetune 85.15 93.95 69.68 69.47 69.07 68.22 8.59
LoRA [20] 4.72 94.07 64.48 62.54 61.63 60.49 6.19
FullLoRA-AT (ours) 4.85 94.04 66.69 66.83 66.22 64.64 6.35
5 Conclusion
@@
In conclusion, to achieve parameter-efficient robustness finetuning for standard pretrained models, our work presents the FullLoRA-AT method, an innovative and lightweight approach designed to significantly enhance the adversarial robustness of pretrained ViT models in a parameter-efficient manner. By integrating the newly proposed LNLoRA modules into several key components of the ViT model, our method can rapidly improve the robustness of the pretrained model with a small number of extra parameters. Extensive experiments show that our method not only addresses critical limitations of previous approaches but also showcases remarkable improvement in the robustness of models against adversarial attacks.
In the future, we plan to apply our method to larger-scale models, such as LLaMA [49], to further validate its effectiveness in more realistic scenarios.

### v2

@@ IV-D Experiments on More Datasets
FullLoRA (ours) 4.85 61.23$\pm$0.12 25.86$\pm$0.05 28.40$\pm$0.04 28.21$\pm$0.03 25.80$\pm$0.13 8.38$\pm$0.02
TinyImageNet Standard Training 84.89 0.00 0.00 0.00 0.00
Fully Finetune 85.37 44.79 24.47 23.46 22.29 22.07 22.96
Aurora [39] 0.09 36.49 18.23 17.92 17.37 17.06 23.49
LoRA [35] 4.72 45.12 21.37 19.79 19.01 18.42 16.29
UniPELT [36] 11.42 43.17 20.56 18.14 17.62 17.04 25.57
FullLoRA (ours) 4.87 45.25$\pm$0.17 23.73$\pm$0.31 22.22$\pm$0.18 21.09$\pm$0.11 20.80$\pm$0.09 16.59$\pm$0.03
ImageNet Standard Training 85.79 0.00 0.00 0.00 0.00
Fully Finetune 86.79 73.82 52.96 51.73 48.29 47.39 207
LoRA [35] 4.86 73.84 47.12 45.61 44.85 42.32 151
FullLoRA (ours) 4.98 73.98 49.37 48.04 46.78 45.17 153
IV-E Results under Different Rank $r$
TABLE V: Results of our FullLoRA under different rank $r$.
@@ IV-E Results under Different Rank $r$
We compare the robustness of pretrained models adversarially finetuned using our proposed FullLoRA method with different ranks $r$ in the LNLoRA module. The experiment is conducted using the ViT-B model on the CIFAR10 dataset. As shown in Table V, as $r$ increases from 8 to 64, clean accuracy improves from 84.52% to 86.82%, and the robustness against various attack methods is also enhanced, with an improvement all over 2%. When comparing the results at $r=64$ with the fully finetuned model, our FullLoRA method utilizes only 11.3% of the training parameters while achieves comparable results and significantly reduces the training time. Considering the trade-off between the number of parameters used for adversarial finetuning and the robustness of the model, we choose rank $r=32$ as the default for other experiments in our work.
@@
TABLE VI: Results of combining FullLoRA with other speed-accelerating adversarial training method AGAT.
Dataset Method Trainable Clean Acc CW-20 PGD-20 PGD-100 AutoAttack Time
Paras (M) (%) (%) (%) (%) (%) (h)
CIFAR-10 AGAT [32] 85.15 87.42 45.02 47.75 47.54 42.86 5.76
FullLoRA (ours) 4.85 87.12 49.76 51.38 50.83 46.94 8.38
CIFAR-100 AGAT [32] 85.22 63.91 25.36 25.74 25.61 24.37 5.77
FullLoRA (ours) 4.85 61.20 25.86 28.40 28.23 25.78 8.38
Imagenette AGAT [32] 85.15 94.70 61.90 61.60 61.10 60.60 4.15
FullLoRA (ours) 4.85 93.40 66.20 66.20 65.60 64.00 6.32
@@
TABLE VII: Results of combining FullLoRA with other speed-accelerating adversarial training methods.
Dataset Method Trainable Clean Acc CW-20 PGD-20 PGD-100 AutoAttack Time
Paras (M) (%) (%) (%) (%) (%) (h)
CIFAR-10 AGAT [32]+LoRA [35] 4.72 86.34 42.68 43.89 43.65 39.68 3.81
AGAT [32]+FullLoRA (ours) 4.85 89.47 44.07 45.00 44.45 41.03 3.93
CIFAR-100 AGAT [32]+LoRA [35] 4.72 65.18 21.73 21.99 21.86 20.42 3.82
AGAT [32]+FullLoRA (ours) 4.85 66.31 22.93 23.07 23.01 22.54 3.93
Imagenette AGAT [32]+LoRA [35] 4.72 93.00 57.20 56.20 56.00 55.10 3.19
AGAT [32]+FullLoRA (ours) 4.85 95.20 60.20 60.00 59.70 58.90 3.23
TABLE VIII: Results of combining FullLoRA with other robustness-enhancing adversarial training methods.
Dataset Method Trainable Clean Acc CW-20 PGD-20 PGD-100 AutoAttack Time
Paras (M) (%) (%) (%) (%) (%) (h)
CIFAR-10 MART [58] 85.15 83.65 51.09 53.62 53.41 49.02 12.13
MART [58]+LoRA [35] 4.72 84.82 47.93 50.77 50.55 46.01 8.86
MART [58]+FullLoRA (ours) 4.85 87.99 50.59 52.09 51.64 47.66 8.97
PRM [61] 85.15 83.43 50.97 53.49 53.12 48.81 11.23
PRM [61]+LoRA [35] 4.72 84.62 47.71 50.53 50.45 45.78 8.16
PRM [61]+FullLoRA (ours) 4.85 87.76 50.39 51.82 51.42 47.36 8.39
CIFAR-100 MART [58] 85.22 61.82 29.51 31.46 31.20 28.23 12.15
MART [58]+LoRA [35] 4.72 61.70 25.44 26.59 26.60 23.36 8.88
MART [58]+FullLoRA (ours) 4.85 62.00 26.60 29.22 29.13 26.68 8.99
PRM [61] 85.22 61.46 29.15 31.12 31.12 28.13 11.25
PRM [61]+LoRA [35] 4.72 62.59 25.11 26.26 26.17 22.97 8.18
PRM [61]+FullLoRA (ours) 4.85 61.81 26.35 28.88 28.70 26.35 8.40
Imagenette MART [58] 85.15 94.26 69.81 69.68 69.40 68.49 9.32
MART [58]+LoRA [35] 4.72 94.06 64.57 62.78 62.00 60.88 6.75
MART [58]+FullLoRA (ours) 4.85 94.19 66.99 67.10 66.51 64.91 6.83
PRM [61] 85.15 93.95 69.68 69.47 69.07 68.22 8.59
PRM [61]+LoRA [35] 4.72 94.07 64.48 62.54 61.63 60.49 6.19
PRM [61]+FullLoRA (ours) 4.85 94.04 66.69 66.83 66.22 64.64 6.35
(a) $\|W_{lora}-W_{adversarial}\|$
(b) $\|W_{fulllore(ours)}-W_{adversarial}\|$
Figure 3: Comparison of the distance between the parameters of different parameter-efficient training methods and the fully finetuned method across the 12 Transformer layers of the ViT model. (a) presents the distance between LoRA and fully finetuned model, and (b) presents the distance between our FullLoRA and fully finetuned model. The experiment is repeated by 5 times, the solid line represents the mean value of the 5 experiments, while the shaded area represents the standard deviation.
IV-F Combined with Other Advanced Adversarial Training Methods
Our method can also be further combined with other advanced adversarial training methods to enhance the results of our adversarial finetuning. The experiment is conducted using the ViT-B model on various datasets.
In Table VI, we present the results of combining FullLoRA with the AGAT method [32], an attention-guided tokens discarding approach designed to accelerate the speed of adversarial training. Comparing our FullLoRA with the AGAT method, it is evident that our FullLoRA achieves significantly higher model robustness than the AGAT method, using only about 5% of the training parameters during adversarial finetuning. Furthermore, in Table VII, we also compare the results of the FullLoRA and LoRA methods when combined with the AGAT method, respectively. It is clear that our FullLoRA also achieves noticeably higher robustness than LoRA, fully demonstrating the superiority of our approach.
Additionally, we also combine our FullLoRA with other advanced methods aimed at enhancing the robustness of adversarial robustness. As shown in Table VIII, when combined with MART [58] and PRM [61], our FullLoRA method still significantly outperforms the LoRA method in terms of model robustness across different datasets.
In summary, when combined with different advanced adversarial training methods, our FullLoRA consistently achieves higher model robustness on various datasets compared to the LoRA method, demonstrating the broad applicability of our proposed FullLoRA method.
IV-G Comparison of the Distance Between Parameters of Different Methods and the Fully Finetuned Method
Compared to LoRA method, our proposed FullLoRA method can more effectively narrow the parameter difference from the fully adversarial training model. To validate it, we conduct experiments to compare the parameter differences between models obtained through various parameter-efficient training methods (e.g., LoRA [35] and our proposed FullLoRA) and models that have been fully finetuned. We denote the model obtained by finetuning a pre-trained model using LoRA as $W_{lora}$, that using our proposed FullLoRA as $W_{ours}$, and that from fully adversarial training as $W_{adversarial}$. We integrate the parameters of the LoRA module into the weights of the Key, Query, and Value variables in the attention module, and the parameters of the layernorm layer in the LNLoRA module are merged into the layernorm layers preceding the MSA and MLP blocks. We visualize $\|W_{lora}-W_{adversarial}\|$ and $\|W_{ours}-W_{adversarial}\|$ of different parameters across the 12 Transformer layers of the ViT model. As shown in Fig. 3, the visualizations reveal that the LoRA method does not effectively reduce the disparity between the pre-trained model and the adversarial model, whereas our proposed method more closely approximates the fully adversarially finetuned model, which well supports our claim that parameter disparities can not be adequately addressed by the existing LoRA method and our proposed LNLoRA aims to learn the differences in feature magnitude between standard and adversarial trained models.
V Conclusion
@@
In conclusion, to achieve parameter-efficient robustness finetuning for standard training models, our work presents the FullLoRA method, an innovative and lightweight approach designed to significantly enhance the adversarial robustness of pretrained ViT models in a parameter-efficient manner. By integrating the newly proposed LNLoRA modules into several key components of the ViT model, our method can rapidly improve the robustness of the pretrained model with a small number of extra parameters. Extensive experiments show that our method not only addresses critical limitations of previous approaches but also showcases remarkable improvement in the robustness of models against adversarial attacks.